OpenStudy (alekos):

find the integral of ln(1/x -x) show all the steps

Parth (parthkohli):

\[\ln\left(\frac{1-x^2}{x}\right)\]\[= \ln(1 +x) + \ln(1-x) - \ln(x)\]Integrate term-by-term.

OpenStudy (anonymous):

Parth that's completely wrong

OpenStudy (anonymous):

That's not right.

OpenStudy (anonymous):

The way to solve this problem is to use the logarithmic integral identity and with substitution.

Parth (parthkohli):

Am I being pranked? :|

OpenStudy (alekos):

i like parthkohli's idea

OpenStudy (alekos):

looks ok to me

OpenStudy (anonymous):

xD sry parth, I'm kidding

OpenStudy (alekos):

thanks Parth. that'll work